Show that the points (2,0,1),(0,4,-3) and
(-2,5,0) are non-collinear.Hence find the equation of plane passing through them.
1
Expert's answer
2020-06-29T19:21:07-0400

three points in three-dimensional space. If they lie on one straight line, then the determinant of the matrix equals 0

"\begin{vmatrix}
 2& 0&1 \\
 0& 4&-3\\
 -2 &5&0
\end{vmatrix}"

the determinant is not equal to 0

therefore the points do not lie on one straight line

finding out the plane passing through three points

"\begin{vmatrix}
 x-x1 & y-y1& z-z1 \\
 x2-x1& y2-y1 &z2-z1\\
 x3-x1 & y3-y1 &z3-z1
\end{vmatrix}" ="\begin{vmatrix}
 x-2 & y& z-1\\
 0-2 & 4-0&-3-1 \\
 -2-2&5 &0-1
\end{vmatrix}" =


= (x-2)*(4*(-1)-(-4)*5)-y*(-2*(-1)-(-4*(-4))+(z-1)*(-2*5-4*(-4))=(x-2)*16-y*(-14)+(z-1)*6=

=16x-32+18y+6z-6=0

got the plane equation

8x+7y+3z=19